Ask your own question, for FREE!
Mathematics 12 Online
OpenStudy (anonymous):

what is the probability of pulling at least 1 ace from a 52 card deck

OpenStudy (perl):

what do you mean by 'at least'. are you drawing 5 cards?

OpenStudy (anonymous):

oh sorry im pulling 3

OpenStudy (perl):

ok. We can solve this using the complement method. P ( at least one ace in three draws ) = 1 - P( no ace in three draws )

OpenStudy (perl):

how many 'non ace' cards are there?

OpenStudy (perl):

$$ \Large { P(A') = 1 - P(A) \\ \therefore \\ P(A) = 1 - P(A~' ) } $$

OpenStudy (anonymous):

thanks!!

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
23 hours ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
1 day ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!